# Careful finite element simulations of cold rolling with accurate through-thickness resolution and prediction of residual stress

### v1

@@ 3.4 Contact conditions
The roller, as an analytically rigid geometry, is defined as a master surface for contact between the roller and the sheet. The surface of the sheet in contact with the roller is defined as the slave (i.e. deformable) surface. Hard pressure-overclosure contact is defined between the roller and the sheet [15]. For cold-rolling processes we expect significant relative motion between the rigid roller and the deformable sheet. To account for relative motion between the roller and sheet surfaces, a finite sliding formulation is adopted. Friction between the roller and the sheet is defined through the isotropic Coulomb friction model. Coulomb friction is regularly used in FE simulations of rolling (e.g., [24, 26, 33, 37, 43]). For industrial cold rolling mills the friction coefficient typically lies in the range of $\mu=0.02-0.15$ [50]. FE models in this work enforce a friction coefficient of $\mu=0.1$.
@@ 3.5 Transient effects and job precision
Our focus in this paper is on steady-state behaviour, so we aim to minimize transient effects during the dynamic explicit analysis.
@@ 3.5.1 Implicit analyses
To ensure computational precision full nodal output was requested for all simulations in this work. This is equivalent to double precision in explicit simulations.
@@ 4 FE model refinement
Here we investigate the effect of element type, solver types (i.e., implicit and explicit) and mesh size on through-thickness stress and strain. Through careful consideration of these variables we obtain a reliable FE model for the rolling process. We then describe the process of ensuring that the simulated rolling process is in a steady state.
@@ 4.1 Element type
In this work, first-order plane-strain elements are employed to discretise the sheet. These elements provide sufficient accuracy and are less computationally-expensive compared to second-order elements [51]. Others [16, 34, 52, 38, 44] similarly used first-order elements to model rolling. As previously discussed (see Section 3.3), ABAQUS limits the element type depending on solver implementation. The element types applied in this work are outlined in Table 3 which differentiates by solver implementation. In the following sections we discuss the benefits (and potential drawbacks) of each element type to demonstrate the effect of element selection on FE model output.
@@ 4.1.1 Implicit analyses
The material response is evaluated by Gaussian quadrature at each element integration point [15]. Plane strain four-node (CPE4) elements contain a total of four integration points. The number of nodes and/or number of integration points correlates to the degree of freedom any given element can exhibit. Reduced integration elements, such as CPE4R elements, use just a single integration point in the centre of the element (compared to the four integration points in a fully integrated element). Therefore, in principle, CPE4 elements are more accurate than the CPE4R elements. However, in the presence of bending the inherent linear geometry of CPE4 elements can give rise to large, artificial shear strains. This problematic behaviour is often termed “shear locking”  [53, 54]. Shear locking is a purely computational behaviour and is not representative of physical processes. FE models exhibiting shear locking provide erroneous displacement and stress values [53].
Figure 2: Shear stress evaluated at the integration points (i.e., no nodal averaging) for a rolling simulation using CPE4 elements. There are significant jumps in shear stress across various elements, creating a checkerboard pattern, which is typical of shear locking behaviour.
Figure 2 shows shear stress evaluated at the integration points of CPE4 elements during a rolling simulation conducted as outlined in the previous sections. In Figure 2 a clear checkerboard pattern is observed, with more extreme shear stress values observed at the corners of elements. This is typical of shear locking behaviour [15] and implies that CPE4 elements are not suitable for this test case, hence this element type is not considered for the remainder of this work.
@@
Unlike CPE4 elements, reduced integration (CPE4R) and “incompatible” (CPE4I) elements do not suffer from shear-locking behaviour. Low-order elements lack the necessary shape functions to describe bending. Incompatible elements (CPE4I) artificially add the required shape functions to adequately describe this behaviour [55, 56]. Simulations conducted using CPE4R and CPE4I elements do not exhibit the checkerboard pattern observed in Figure 2. This indicates that both reduced-integration and incompatible elements are sufficient for removing shear locking behaviour. However, incompatible elements are slow to convergence in simulations with large compressive strains [15]. This severely restricts the mesh size that can be employed with the CPE4I element type (see Section 3 of the supplementary information provided, where the performance of CPE4I and CPE4R elements are compared), and hence CPE4I elements are not considered for the remainder of this work.
@@ 4.2 Limitations of ABAQUS/Explicit
We now consider results from ABAQUS/Explicit simulations. Mass scaling (or increased loading rate) was not used in any explicit simulation in this paper. Given that the kinetic-to-internal energy ratio remains in the desired range (5-10%), we consider inertia effects to be insignificant.
Figure 3: Contour plots showing von Mises stress ($\mathrm{M}\mathrm{P}\mathrm{a}$) from explicit simulations with (a) $N_{e}=5$ CPE4R elements and (b) $N_{e}=40$ CPE4R elements. Both simulations have roller radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
As mesh density increases, it is expected that the calculation accuracy should, in general, improve [61]. However, Figure 3 shows that increasing the mesh density increases noise. Figure 3 compares the von Mises stress for two explicit simulations with mesh densities of $N_{e}=5$ (Figure 3a) and $N_{e}=40$ (Figure 3b) elements through the half-thickness, respectively. The most notable difference between the plots is the smoothness of the solution. This result indicates that the solution diverges as mesh density increases. In particular, we note the noisy oscillatory behaviour near the surface of the sheet towards the entrance to the roll gap, $x/L=0$. Flanagan et al. [11] successfully removed noise by implementing additional stiffness-proportional damping. The damping approach however increases the required computational time [700 CPU hours compared to 10 CPU hours without damping for a particular rolling simulation conducted by 11]. Given the divergent nature of the explicit results and considering the lack of equilibrium checks we propose that implicit simulations outperform explicit simulations for this case study. Therefore the remainder of this work will focus on simulation results from implicit simulations only.
@@ 4.3 Mesh sensitivity study
In subsection 4.1 we demonstrated that CPE4R elements are the preferred element type for this simulation case study. We also demonstrated (subsection 4.2) that explicit analyses are subject to increasing noise with increasing mesh density. In this section we perform a convergence study for implicit simulations that results in the optimum trade-off between accuracy and computational intensity.
Figure 4: Time-history plots of (a) roll force and (b) torque per unit width during the rolling analysis step for implicit simulations with $N_{e}=5$, $N_{e}=30$ and $N_{e}=40$ CPE4R elements. Relatively steady average force and torque values are observed after $0.001\,$\mathrm{s}$$ for each simulation. All simulations have radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
Figure 4 plots (a) roll force and (b) torque per unit width as a function of simulation time during the rolling step for various mesh densities. We note that the $N_{e}=5$ simulation gives average roll force and torque values similar to the values outputted from the simulations with $N_{e}=30$ and $N_{e}=40$. Some authors use roll force and torque in their mesh convergence studies by identifying the mesh size at which no further significant changes in force or torque occur [62, 18]. By extension, we could take the $N_{e}=5$ results as sufficient here. However, as we will see in Figure 5, the simulation with $N_{e}=5$ performs poorly in predicting the von Mises stress, shear stress and plastic equivalent strain (PEEQ) values through the thickness of the sheet, thus proving that roll force and torque are poor measures of accuracy when through-thickness predictions are desired.
@@
For von Mises stress, the largest and smallest numerical errors occur near $x/L=1.09$ (see the dashed lines in Figure 5(a)) and $x/L=0.6913$ (see the dotted lines in Figure 5(a)) respectively. This makes sense because the von Mises stress satisfies the yield condition uniformly in the interior of the contact region, except very close to the start and end where the material may be behaving elastically. For the shear stress, the largest errors occur at $x/L=0.025$ (see the solid curves in Figure 5(b)). The maximum relative difference in shear between the simulations with $N_{e}=30$ and $N_{e}=40$ is 4.9% at this position. Finally, the deviations in PEEQ between different mesh densities are largest near $x/L=1.09$ (see Figure 5(c)), but the PEEQ errors are much smaller than for von Mises stress and shear stress (see Table 4).
We note that the simulation with $N_{e}=5$ does not capture trends such as non-monotonic von Mises stress at $x/L=1.09$ (Figure 5(a)) or the curvature of the shear stress (Figure 5(b)). Yarita et al. [28] and Cawthorn et al. [39] used just $N_{e}=4$ and $N_{e}=5$ reduced-integration elements respectively through the half-thickness of the sheet, which our results suggest is far too coarse a mesh to accurately capture the through-thickness stress and strain distributions in the sheet. We use $N_{e}=30$ CPE4R elements through the half-thickness of the sheet in the remainder of this work since the maximum difference from the simulation with $N_{e}=40$ is 4.9%, and the computation time is 4.83 times faster compared to the $N_{e}=40$ simulation.
Figure 5: Interpolated (a) von Mises stress, (b) shear stress and (c) PEEQ (plastic equivalent strain) from simulations with varying mesh densities. The solid, dotted and dashed lines depict results in the sheet at $x/L=0.025$, $x/L=0.6913$ and $x/L=1.09$, respectively (see Figure 1). The horizontal centre (symmetry axis in Figure 1) of the sheet corresponds to $z/h_{0}=0$. All simulations employ a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
@@ 4.4 Steady-state determination
Figure 6 plots interpolated von Mises stress, shear stress and PEEQ results from different time points throughout the $N_{e}=30$ simulation. Although Figure 4 suggests that the roll force and torque results reach a relatively steady value after time $t=0.001\,$\mathrm{s}$$, the interpolated results in Figure 6 suggest that the through-thickness stress and strain components are still significantly changing in time. This highlights the inadequacy of roll force and torque in terms of predicting the steady state of the rolling process. Despite this inadequate steady-state determination, roll force and torque are used by various authors [see, for example, 16, 43] for determining if a steady state has been reached by analysing the time at which these quantities become relatively constant. Although the roll force and torque are necessarily constant during the steady state of the rolling process, constant roll force and torque are not sufficient for identifying steady-state behaviour on their own, as shown here.
Figure 6 suggests the system is in a steady state by $t=0.0802\,$\mathrm{s}$$, since the through-thickness predictions do not significantly change from this time onwards. The data used for the remainder of this paper is therefore extracted at $t=0.0802\,$\mathrm{s}$$ to both ensure steady results and to facilitate the velocity calculation in the postprocessing (see Section 2 of the supplementary material).
@@
Figure 6: Interpolated (a) von Mises stress, (b) shear stress and (c) PEEQ (plastic equivalent strain) from different time points of the implicit simulation with $N_{e}=30$ CPE4R elements. The solid, dotted and dashed lines depict results in the sheet at $x/L=0.025$, $x/L=0.6913$ and $x/L=1.09$, respectively (see Figure 1). The horizontal centre (symmetry axis in Figure 1) of the sheet corresponds to $z/h_{0}=0$. The simulation employs a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
5 Comparison to analytical models

### v2

@@ 3.4 Contact conditions
The roller, as an analytically rigid geometry, is defined as the master surface for contact between the roller and the sheet. The surface of the sheet in contact with the roller is defined as the slave (i.e. deformable) surface. Hard pressure-overclosure contact is the most common contact pressure-overclosure relationship and is defined between the roller and the sheet [22]. In theory, by using this hard contact relationship, contact pressure is zero until contact actually occurs, and immediately increases to infinity upon contact. This infinite contact force prevents penetration of the slave surface into the master surface. In practice, however, this infinite spike in contact pressure is numerically challenging. The zero-penetration condition may or may not be strictly enforced depending on the constraint enforcement method used. The default penalty constraint enforcement method is implemented here, which approximates hard pressure-overclosure behavior, so some degree of penetration will occur. For cold-rolling processes we expect significant relative motion between the rigid roller and the deformable sheet. To account for relative motion between the roller and sheet surfaces, a finite sliding formulation is adopted. Friction between the roller and the sheet is defined through the isotropic Coulomb friction model. Coulomb friction is regularly used in FE simulations of rolling (e.g., [40, 9, 12, 15, 18]). For industrial cold rolling mills the friction coefficient typically lies in the range of $\mu=0.02-0.15$ [55]. FE models in this work enforce a friction coefficient of $\mu=0.1$.
@@ 3.5 Transient effects and job precision
Our focus in this paper is on steady-state behaviour, so we aim to minimise transient effects during the dynamic explicit analysis as outlined in subsection 3.5.2.
@@ 3.5.1 Implicit analyses
To ensure computational precision, full nodal output was requested for all simulations in this work. This is equivalent to double precision in explicit simulations.
@@ 4 FE model refinement
Here we investigate the effect of element type, solver types (i.e., implicit and explicit) and mesh size on through-thickness stress and strain predictions. Through careful consideration of these variables we obtain a reliable FE model for the rolling process. We then describe the process of ensuring that the simulated rolling process is in a steady state.
@@ 4.1 Element type
In this work, first-order plane-strain elements are employed to discretise the sheet. These elements provide sufficient accuracy and are less computationally-expensive compared to second-order elements [56]. Others [23, 45, 57, 46, 35] similarly used first-order elements to model rolling. As previously discussed (see subsection 3.3), ABAQUS limits the element type depending on solver implementation. The element types applied in this work are outlined in Table 3 which differentiates by solver implementation. In the following sections we discuss the benefits (and potential drawbacks) of each element type to demonstrate the effect of element selection on FE model output.
@@ 4.1.1 Implicit analyses
The material response is evaluated by Gaussian quadrature at each element integration point [22]. Fully-integrated plane-strain four-node (CPE4) elements contain a total of four integration points. The number of nodes and/or number of integration points correlates to the degree of freedom any given element can exhibit. Reduced-integration elements, such as CPE4R elements, use just a single integration point in the centre of the element (compared to the four integration points in a fully-integrated CPE4 element). Therefore, in principle, CPE4 elements are more accurate than the CPE4R elements. However, in the presence of bending moments the inherent linear geometry of CPE4 elements can give rise to large, artificial shear strains. This problematic behaviour is often termed “shear locking” [58, 59]. Shear locking is a purely computational behaviour and is not representative of physical processes. FE models exhibiting shear locking provide erroneous displacement and stress values [58].
Figure 3: Interpolation is carried out in Figures 4, 7 and 8 in the $z$-direction to facilitate comparison between different FE results. The solid, dotted and dashed curves in those figures depict results in the sheet at $x=0.025\,L$, $x=0.6913\,L$ and $x=1.09\,L$, respectively, as shown here. These are the three $x$-positions of the maximum relative errors (see D), which are calculated with respect to the simulation with $N_{e}=40$.
Figure 4: Interpolated shear results from two separate simulations that employ $N_{e}=40$ CPE4 and CPE4R elements through the half-thickness respectively. Readers should refer to Figure 3 which identifies the horizontal position for the solid, dotted and dashed lines, respectively. The horizontal centre (symmetry axis in Figure 1) of the sheet corresponds to $z/h_{0}=0$. Both simulations employ a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and $\varepsilon=0.125$.
Figure 14 shows a checkerboard pattern in the shear stress which is evaluated at the integration points of CPE4 elements during a rolling simulation with $N_{e}=40$ elements through the half-thickness of the sheet. This checkerboard pattern is typical of shear locking behaviour [22]. Figure 4 shows a quantitative comparison between the aforementioned CPE4 shear results and shear outputs from a simulation that employs $N_{e}=40$ CPE4R elements through the half-thickness. Interpolation is carried out in the $z$-direction to facilitate comparison between different mesh types. The solid, dotted and dashed curves depict results in the sheet at $x=0.025\,L$, $x=0.6913\,L$ and $x=1.09\,L$, respectively (see Figure 3). These three $x$-positions represent the positions where maximum relative errors were observed in our convergence analysis (see D for more details). Figure 4 shows sharp changes in the nodal CPE4 shear results, particularly towards the surface of the sheet (i.e., higher $z/h_{0}$ values) at $x=0.025\,L$ and $x=0.6913\,L$ (the solid and dotted curves respectively), highlighting the non-smooth nature of the shear results that is also visible in Figure 14. In contrast, the CPE4R results are quite smooth everywhere, confirming that shear locking is not an issue for CPE4R elements. The checkerboard pattern (Figure 14) and the interpolated shear results (Figure 4) imply that CPE4 elements are not suitable for this test case and hence this element type is not considered for the remainder of this work.
@@
Unlike CPE4 elements, plane-strain reduced-integration (CPE4R) and incompatible (CPE4I) elements do not suffer from shear-locking behaviour. Low-order elements lack the necessary shape functions to describe bending. Incompatible elements such as CPE4I elements artificially add the required shape functions to adequately describe this behaviour [60, 61]. Simulations conducted using CPE4R and CPE4I elements do not exhibit the typical shear-locking checkerboard pattern observed in Figure 14. This indicates that employing either reduced-integration or incompatible elements would be sufficient for removing shear-locking behaviour. However, incompatible elements are slow to converge in simulations with large compressive strains [22]. For example, a simulation with $N_{e}=15$ CPE4I elements through the half-thickness took over 5.5 times as long as a simulation with $N_{e}=15$ CPE4R elements, (see C, where the performance of CPE4I and CPE4R elements are compared in more detail). The large computational time severely restricts the mesh size that can be employed with the CPE4I element type, and hence CPE4I elements are not considered for the remainder of this work.
@@ 4.2 Limitations of ABAQUS/Explicit
We now consider results from ABAQUS/Explicit simulations. Mass scaling (or increased loading rate) was not used in any explicit simulation in this paper. Given that the kinetic-to-internal energy ratio remains in the desired range (5–10%) [67, 22], we consider inertia effects to be insignificant.
Figure 5: Contour plots showing von Mises stress ($\mathrm{M}\mathrm{P}\mathrm{a}$) from explicit simulations with (a) $N_{e}=5$ CPE4R elements and (b) $N_{e}=40$ CPE4R elements, where the symmetry axis is located at $z/h_{0}=0$ (see Figure 1). Both simulations have roller radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and $\varepsilon=0.125$.
Figures 5(a) and 5(b) compare the von Mises stress for two explicit simulations with mesh densities of $N_{e}=5$ and $N_{e}=40$ elements through the half-thickness, respectively. The most notable difference between the plots is the smoothness of the solution. As mesh density increases, it is expected that the calculation accuracy should, in general, improve [68]. However, Figure 5 shows that increasing the mesh density increases noise, indicating that the solution diverges as mesh density increases. In particular, we note the numerical noise near the surface of the sheet towards the entrance to the roll gap, $x/L=0$. This noise is likely an artefact of the lack of convergence checks performed by ABAQUS/Explicit at each time increment, which allows the solution to diverge without restriction. Flanagan et al. [32] successfully removed such noise by implementing additional stiffness-proportional damping. However, the damping approach increases the required computational time (700 CPU hours compared to 10 CPU hours without damping for a particular rolling simulation conducted by Flanagan et al. [32]). Although the explicit simulations are typically quicker to execute than their implicit counterparts (for example, with $N_{e}=40$ CPE4R elements through the half-thickness, the explicit simulation’s CPU time is less than 37% than that of the equivalent implicit simulation), the divergent nature of the explicit results and the lack of equilibrium checks lead us to proposing that implicit simulations outperform explicit simulations for this case study. Therefore the remainder of this work will focus on simulation results from implicit simulations only.
@@ 4.3 Mesh sensitivity study
In subsection 4.1 we demonstrated that CPE4R elements are the preferred element type for this simulation case study. We also demonstrated (subsection 4.2) that explicit analyses are subject to increasing noise with increasing mesh density. In this section we perform a mesh convergence study for implicit simulations with CPE4R elements that results in the optimum trade-off between accuracy and computational intensity.
Figure 6: Time-history plots of (a) roll force and (b) torque per unit width during the rolling analysis step for implicit simulations with $N_{e}=5$, $N_{e}=30$ and $N_{e}=40$ CPE4R elements. Relatively steady average force and torque values are observed after $0.001\,$\mathrm{s}$$ for each simulation. All simulations have radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and $\varepsilon=0.125$.
@@
[…]
We note that the simulation with $N_{e}=5$ does not capture trends such as non-monotonic von Mises stress at $x=1.09\,L$ (Figure 7(a)) or the curvature of the shear stress (Figure 7(b)). Yarita et al. [10] and Cawthorn et al. [17] used just $N_{e}=4$ and $N_{e}=5$ reduced-integration elements respectively through the half-thickness of the sheet, which our results suggest is far too coarse a mesh to accurately capture the through-thickness stress and strain distributions in the sheet. We use $N_{e}=30$ CPE4R elements through the half-thickness of the sheet in the remainder of this work since the maximum difference from the simulation with $N_{e}=40$ is 4.9%, and the computation time is 4.83 times faster compared to the $N_{e}=40$ simulation.
Figure 7: Comparing the number of elements through half-thickness $\left(N_{e}\right)$ for simulations with radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and $\varepsilon=0.125$. Plotting interpolated (a) von Mises stress, (b) shear stress and (c) PEEQ (plastic equivalent strain). Readers should refer to Figure 3 which identifies the horizontal position for the solid, dotted and dashed lines, respectively.
@@ 4.4 Steady-state determination
In Section 2 we noted that steady-state conditions are often defined in FE simulations through roll-force and/or roll-torque results in the literature. Figure 6 suggests that for all mesh sizes, both roll force and torque reach a relatively steady value by time $t\approx 0.001\,$\mathrm{s}$$. With the help of Figure 8 however we demonstrate here how roll force and torque are insufficient at truly assessing whether steady-state conditions have been achieved, and show how to accurately determine if a true steady state has been reached.
@@
Figure 8: Comparing results from different time points of the implicit simulation with $N_{e}=30$ CPE4R elements, radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and $\varepsilon=0.125$, by plotting interpolated (a) von Mises stress, (b) shear stress and (c) PEEQ (plastic equivalent strain). Readers should refer to Figure 3 which identifies the horizontal position for the solid, dotted and dashed lines, respectively.
Figure 8 shows interpolated von Mises stress, shear stress and PEEQ results from six different time points throughout the $N_{e}=30$ simulation. Interpolation is carried out as outlined in subsection 4.3. The purpose of Figure 8 is to assess the time at which a true through-thickness steady state has been achieved. Since each set of curves (i.e., solid, dotted and dashed) represent the interpolated results at a particular $x$-position, we expect that under steady-state conditions the shape of each set of through-thickness curves should remain stagnant as a function of time. That is to say that, in a steady state, the profile of the curve should not change from time-to-time because significant differences from consecutive time increments indicates that a steady state has not been achieved.
Firstly we draw the readers attention to Figure 8(a) which details the von Mises stress as a function of through-thickness position ($z/h_{0}$) for the simulation with $N_{e}=30$ CPE4R elements through the half-thickness. We previously showed this element size and type to provide the optimal trade off between computational expense and accuracy. In Figure 8(a) we plot results from six time points throughout the simulation, at identical horizontal positions that were analysed previously (see Figure 3). The solid and dotted curves in Figure 8(a) all have very similar profiles to their respective curves for all time points throughout the simulation. However, the dashed curves (representing results from $x=1.09\,L$) reveal unsteady behaviour at times $t=0.0011\,$\mathrm{s}$$ and $t=0.0022\,$\mathrm{s}$$, where these curves are noticeably different to the $t\geq 0.0033\,$\mathrm{s}$$ curves at the same $x$-position. Figure 8(b) depicts the interpolated shear results at the same three $x$-positions. Once again, the solid curves all lie on top of each other, showing no unsteady behaviour at $x=0.025\,L$ in terms of shear stress. However, the dotted and dashed curves vary significantly from time-to-time until $t>0.0033\,$\mathrm{s}$$. Similarly, the PEEQ (Figure 8(c)) results at $x=1.09\,L$ (dashed curves) do not indicate steady-state conditions until time $t>0.0033\,$\mathrm{s}$$. It is therefore clear that the through-thickness stress and strain states do not reach a steady state at all $x$-positions until $t>0.0033\,$\mathrm{s}$$. However, as pointed out earlier, Figure 6 suggests that roll force and torque reach a steady value by $t=0.001\,$\mathrm{s}$$, highlighting the inadequacy of roll force and torque in terms of predicting the steady state of the rolling process when through-thickness accuracy is desired since stress and strain components are still significantly changing at $t=0.001\,$\mathrm{s}$$, as shown here. Despite this inadequate steady-state determination, roll force and torque are used by various authors [see, for example, 23, 18] for determining if a steady state has been reached.
The interpolated results from the last three time points in Figure 8 are indistinguishable for all stress and strain quantities at all three $x$-positions, and so we can say with confidence that a steady state is reached by $t=0.0802\,$\mathrm{s}$$. Hence, we extract the data used for the remainder of this paper at $t=0.0802\,$\mathrm{s}$$ to both ensure steady results and to facilitate the velocity calculation in the postprocessing (see B).
5 Numerical results